In youngs double slit experiment with wavelength 5890A°; there are 60 fringes in the field of vision. How many fringes will be observed in the same field of vision if wavelength used is5460A°

Grade:12

1 Answers

Vikas TU
6 years ago
Number of edges is straightforwardly relative to the wavelength 
 
N1 α ƛ1 
 
N2 α ƛ2 
 
N1/N2 = ƛ1/ƛ2 
 
60/N2 = 5890/5460 
 
N2 = (5460 *60)/5890 
 
N2 = 55